OpenStudy (anonymous):

I will be your fan and give you a medal What is 5/18+(-1/6) A. 1/9 B. 4/9 C. 2/3 D. 1/3 I think it's B

OpenStudy (mathmale):

Here's what I'd suggest you do up front: (1) simplify that +(-1/6) by rewriting it as -1/6 (2) identifying the lowest common denominator of the two fractions (3) writing both fractions with that LCD (4) combining the 2 fractions.

OpenStudy (mathmale):

Hint: does 6 divide into 18?

OpenStudy (anonymous):

Yes it does

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

sorry for the delays. 5/18+(-1/6) can be written as\[\frac{ 5 }{ 18 }-\frac{ 1 }{ 6 }\] If it is, you may find it easier to identify the LCD. You already know that 6 divides into 18 just fine. So 18 is the LCD. How can you "fix" the right hand fraction so that it will have the same denominator as 5/18?

OpenStudy (anonymous):

1/18

OpenStudy (mathmale):

I see you';ve multiplied that 6 in the denom. by 3, to get 18. But, by doing so, you've changed the value of the fraction from 1/6 to 1/18. Realize that if you mult. the den. by 3, you must also mult. the numerator by 3. So, the first fraction is 5/18; the second fraction is what?

OpenStudy (mathmale):

\[\frac{ 1 }{ 6 }*\frac{ 3 }{ 3 }=?\]

OpenStudy (anonymous):

3/6

OpenStudy (mathmale):

Wouldn't that be 1*3 = 3 OVER 6*3 = 18?

OpenStudy (mathmale):

\[\frac{ 5 }{ 18 }-\frac{ 3 }{ 18 }=?\]

OpenStudy (anonymous):

2/18

OpenStudy (anonymous):

THIS is SO confusing..I dont even undestand wht you're trying to write

OpenStudy (mathmale):

Must have the same den. in each fract. before we can add or subtr. the fractions. Good. Now reduce that fraction, please.

OpenStudy (mathmale):

(1) simplify that +(-1/6) by rewriting it as -1/6 (2) identifying the lowest common denominator of the two fractions (3) writing both fractions with that LCD (4) combining the 2 fractions.

OpenStudy (anonymous):

1/9

OpenStudy (mathmale):

that's correct. Very nice. Please review what we were aiming to accomplish, along with the steps of the procedure that I've repeated, above. Rule: If you want to combine 2 fractions, they must have the same den. That's why we've changed 1/6 to 3/18.

OpenStudy (anonymous):

Ok I will review because this is very important!!!

OpenStudy (mathmale):

So, Jada, are you satisfied with what we've done? If you like, you could post another problem involving combining fractions and we could work on that new problem. But please post it separately.

OpenStudy (anonymous):

Ok and yess im satisifed

OpenStudy (mathmale):

Cool. Hope to "meet" you again on OpenStudy! Best wishes to you. MM